# Graphical Calculus for Non-Gaussian Quantum States

### v2

@@ Appendix A Derivations of graphical calculus
Displacement in momentum: The momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ modifies the hyperedge set by $(E,t)\rightarrow(E,t)\cup(\{a\},s)$. That is, it adds the hyperedge $\{a\}$ with the weight $s$ if there was no such hyperedge before or changes the weight by $s$.
@@
Displacement in position: Applying the position displacement operator $X(s)$ on the mode $a$ modifies the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ such that $(E^{\prime},t^{\prime})=(E,t)\cup(\operatorname{\mathcal{A}}(a),\tilde{t})$, where $\tilde{t}_{e}\coloneqq-s\times t_{e\cup a}$. In other words, it adds or updates the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ and multiplies the weight of the hyperedge containing $a$ by $s$:
@@
Shearing: Applying shearing in position is trivial. This expression
@@
One-mode squeezing: Applying a one-mode squeezing operation $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}=\operatorname{e}^{-\frac{s}{2}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}$ on the mode $a$ changes the weight of all hyperedges containing $a$ by a factor of $\operatorname{e}^{-s}$:
@@
Rotation: Applying a rotation operator in general gives us a state which is not in the form of a hypergraph state. The general form is given by:
@@
Controlled-$Z$ gate: The effect of application of the controlled-$Z$ $C_{a,b}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{q}}}_{b}}$ gate on two modes $a,b$ directly follows from the definition of hypergraph states. If the edge $\{a,b\}\in E$ was part of the edge set, its weight gets modified, since
@@
Measurement in $\operatorname{\mathit{\hat{q}}}$ basis: Measuring |$\outerproduct{m}{m}_{q,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$. The hyperedge set and weights get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@
Measurement in $\operatorname{\mathit{\hat{p}}}$ basis: Measuring |$\outerproduct{m}{m}_{p,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$ and creates several hyperedges in a superposition of all possible weights, with a phase determined by the weight and the measurement result $m$. The hyperedge set and weight set get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@ Appendix B Effects of Finite Squeezing
The fidelity of |$\ket{H_{\text{predict}}}$ and the actually displaced state $X^{A}(s)\ket{H_{\text{approx}}}$ is given by $\operatorname{e}^{-\frac{1}{2}s^{2}\operatorname{e}^{-2r_{A}}}$, which tends to 1 for a large squeezing parameter $r_{A}$. Here, we assume that the states can be squeezed such that the fidelity between the actually prepared state and the state computed using the graphical rules from this paper is sufficiently high. Other rules can be calculated fidelity in a similar way.
Figure 5: Generation of a state with $m$ 3rd-order hyperedges. (a) The initial hypergraph state with only one 3rd-order hyperedge. (b) After performing momentum shearing on mode $A$, the desired purple hyperedges of $\{B,C,D_{1}\}$,…,$\{B,C,D_{m}\}$ are generated. Additionally, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{1}}^{2}}$,…, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{m}}^{2}}$, and $e^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}}$, are produced as by-products. (c) Removing the $\{A,D_{1}\}$, …, $\{A,D_{m}\}$ edge by performing a series of two-mode controlled-$Z$ gates, $C_{A,D_{1}}(-1)$, …, $C_{A,D_{m}}(-1)$. (d) After performing the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}}$$\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.
@@
In Fig. 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state.
We can see that, after performing the shearing operations on mode $A$ and a series of controlled-$Z$ gates as well as the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state Instead of applying a non-Gaussian Toffoli gate on a certain hypergraph state, we can apply a sequence of Gaussian operations on the same state which has the same effect.

### v3

@@ Appendix A Derivations of the graphical framework
A.1 Displacement in momentum
The momentum displacement operator $Z^{a}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}}$ modifies the hyperedge set by $(E;t)\rightarrow(E;t)\cup(\{a\};s)$. That is, it adds the hyperedge $\{a\}$ with the weight $s$ if there was no such hyperedge before or changes the weight by $s$.
@@ A.1 Displacement in momentum
A.2 Displacement in position
Applying the position displacement operator $X(s)$ on the mode $a$ modifies the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ such that $(E^{\prime};t^{\prime})=(E;t)\cup(\operatorname{\mathcal{A}}(a);\tilde{t})$, where $\tilde{t}_{e}\coloneqq-s\times t_{e\cup a}$. In other words, it adds or updates the hyperedges in the adjacency $\operatorname{\mathcal{A}}(a)$ of $a$ and multiplies the weight of the hyperedge containing $a$ by $s$:
@@ A.2 Displacement in position
A.3 Shearing
Applying shearing in position is trivial. This expression
@@ A.3 Shearing
A.4 One-mode squeezing
Applying a one-mode squeezing operation $S^{a}(s)=\operatorname{e}^{-i\frac{s}{2}(\operatorname{\mathit{\hat{p}}}_{a}\operatorname{\mathit{\hat{q}}}_{a}+\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a})}=\operatorname{e}^{-\frac{s}{2}}\operatorname{e}^{-is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{p}}}_{a}}$ on the mode $a$ changes the weight of all hyperedges containing $a$ by a factor of $\operatorname{e}^{-s}$:
@@ A.4 One-mode squeezing
A.5 Rotation
Applying a rotation operator in general gives us a state which is not in the form of a hypergraph state. The general form is given by:
@@ A.5 Rotation
The rotation operator can be decomposed into a sequence of squeezing and shearing operators. In the following, we show that $R(s)=D_{p}(t)S(r)D_{q}(t)$ for $t=\tan(s)$ and $e^{r}=1/\cos(s)$ for rotations $s\neq\frac{(2n+1)\pi}{2}$, where $n\in\operatorname{\mathds{N}}$. The rotation, squeezing, and shearing operators in the phase space notation are given by
$D_{q}(t)$ $=\begin{pmatrix}1&0\\ t&1\end{pmatrix},\quad$ $S(r)$ $=\begin{pmatrix}e^{r}&0\\ 0&e^{-r}\end{pmatrix},$
$D_{p}(t)$ $=\begin{pmatrix}1&-t\\ 0&1\end{pmatrix},\quad$ $R(s)$ $=\begin{pmatrix}\cos(s)&-\sin(s)\\ \sin(s)&\cos(s)\end{pmatrix}.$ (12)
It can be checked that $R(s)=D_{p}(t)S(r)D_{q}(t)$ holds for $t=\tan(s)$ and $e^{r}=1/\cos(s)$.
A.6 Controlled-$Z$ gate
The effect of application of the controlled-$Z$ $C_{a,b}(s)=\operatorname{e}^{is\operatorname{\mathit{\hat{q}}}_{a}\operatorname{\mathit{\hat{q}}}_{b}}$ gate on two modes $a,b$ directly follows from the definition of hypergraph states. If the edge $\{a,b\}\in E$ was part of the edge set, its weight gets modified, since
@@ A.6 Controlled-$Z$ gate
A.7 General Gaussian Operations
As discussed in the main text, general Gaussian operations can be decomposed into a sequence of the discussed operations [54, 50, 45, 55]. In addition to the decompositions mentioned in the literature, we prove the decomposition of the rotation operator in terms of squeezing and shearing operators in Sect. A.5.
Not all the discussed operators map hypergraph states to hypergraph states. The operations shearing and rotation map hypergaph states to hypergraph states up to remaining unitaries. The remaining unitary of a hypergraph state associated to the hypergraph $H=(V,E)$ are of the form $e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$, where $f(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})$ is a polynomial function of momentum and position operators of the modes corresponding to the vertex set $V$. In this section we show how the remaining unitaries in the most general form commute with the operators. We explicitly derive the relation for the shearing operator $D_{q}(s)$, the other relations are derived equivalently.
For computing the commutation relation of the shearing operator $D_{q}(s)$, we use the identities
$D_{q}(s)\operatorname{\mathit{\hat{q}}}D_{q}(-s)$ $=\operatorname{\mathit{\hat{q}}},$ $D_{q}(s)\operatorname{\mathit{\hat{q}}}^{n}D_{q}(-s)$ $=(\operatorname{\mathit{\hat{q}}})^{n},$
$D_{q}(s)\operatorname{\mathit{\hat{p}}}D_{q}(-s)$ $=\operatorname{\mathit{\hat{p}}}-s\operatorname{\mathit{\hat{q}}},$ $D_{q}(s)\operatorname{\mathit{\hat{p}}}^{n}D_{q}(-s)$ $=(\operatorname{\mathit{\hat{p}}}-s\operatorname{\mathit{\hat{q}}})^{n}.$ (14)
Furthermore, we have
$D_{q}(s)f(\operatorname{\mathit{\hat{p}}},\operatorname{\mathit{\hat{q}}})D_{q}(-s)=f(D_{q}(s)\operatorname{\mathit{\hat{p}}}D_{q}(-s),D_{q}(s)\operatorname{\mathit{\hat{q}}}D_{q}(-s)).$ (15)
This can be seen by inserting identities $\operatorname{\mathds{1}}=D_{q}(-s)D_{q}(s)$ between all products of operators in the function. Using the Taylor expansion, we get
$D_{q}(s)e^{if(\operatorname{\mathit{\hat{p}}},\operatorname{\mathit{\hat{q}}})}D_{q}(-s)=e^{if(D_{q}(s)\operatorname{\mathit{\hat{p}}}D_{q}(-s),D_{q}(s)\operatorname{\mathit{\hat{q}}}D_{q}(-s))}.$ (16)
For a multimode remaining unitary, we therefore have
$D_{q}^{a}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}=e^{if(p_{1},q_{1},...p_{a}-sq_{a},q_{a},...p_{\lvert V\rvert},q_{\lvert V\rvert})}D_{q}^{a}(s).$ (17)
The commutation relations of other operators from Sect. 3 and the remaining unitary term are listed below.
$X^{a}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$ $=e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{a},\operatorname{\mathit{\hat{q}}}_{a}+s,\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}X^{a}(s),$
$Z^{a}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$ $=e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{a}-s,\operatorname{\mathit{\hat{q}}}_{a},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}Z^{a}(s),$
$D_{p}^{a}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$ $=e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{a},\operatorname{\mathit{\hat{q}}}_{a}+s\operatorname{\mathit{\hat{p}}}_{a},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}D_{p}^{a}(s),$
$D_{q}^{a}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$ $=e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{a}-s\operatorname{\mathit{\hat{q}}}_{a},\operatorname{\mathit{\hat{q}}}_{a},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}D_{q}^{a}(s),$
$S^{a}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$ $=e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,e^{-s}\operatorname{\mathit{\hat{p}}}_{a},e^{s}\operatorname{\mathit{\hat{q}}}_{a},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}S^{a}(s),$
$C_{a,b}(s)e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}$ $=e^{if(\operatorname{\mathit{\hat{p}}}_{1},\operatorname{\mathit{\hat{q}}}_{1},\dots,\operatorname{\mathit{\hat{p}}}_{a}-\operatorname{\mathit{\hat{q}}}_{b},\dots,\operatorname{\mathit{\hat{p}}}_{b}-\operatorname{\mathit{\hat{q}}}_{a},\dots,\operatorname{\mathit{\hat{p}}}_{\lvert V\rvert},\operatorname{\mathit{\hat{q}}}_{\lvert V\rvert})}C_{a,b}(s).$ (18)
By using these identities, we can track the propagated remaining unitaries and defer them to the end.
A.8 Measurement in $\operatorname{\mathit{\hat{q}}}$ basis
@@
Measuring |$\outerproduct{m}{m}_{q,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$. The hyperedge set and weights get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@
A.9 Measurement in $\operatorname{\mathit{\hat{p}}}$ basis
Measuring |$\outerproduct{m}{m}_{p,a}$ disconnects mode $a$ from the hypergraph by disconnecting hyperedges containing $a$ and creates several hyperedges in a superposition of all possible weights, with a phase determined by the weight and the measurement result $m$. The hyperedge set and weight set get updated as $E^{\prime}=(E\setminus E_{a})\cup\operatorname{\mathcal{A}}(a)$ and for $e\in E^{\prime}$
@@ Appendix B Effects of Finite Squeezing
The fidelity of |$\ket{H_{\text{predict}}}$ and the actually displaced state $X^{A}(s)\ket{H_{\text{approx}}}$ is given by $\operatorname{e}^{-\frac{1}{2}s^{2}\operatorname{e}^{-2r_{A}}}$, which tends to 1 for a large squeezing parameter $r_{A}$. Here, we assume that the states can be squeezed such that the fidelity between the actually prepared state and the state computed using the graphical rules from this paper is sufficiently high. The fidelity for other operations can be computed equivalently.
@@
Figure 5: Generation of a state with $m$ 3rd-order hyperedges. (a) The initial hypergraph state with only one 3rd-order hyperedge. (b) After performing momentum shearing on mode $A$, the desired purple hyperedges of $\{B,C,D_{1}\}$,…,$\{B,C,D_{m}\}$ are generated. Additionally, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{1}}^{2}}$,…, $\operatorname{e}^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{D_{m}}^{2}}$, and $e^{i\frac{1}{2}\operatorname{\mathit{\hat{q}}}_{B}^{2}\operatorname{\mathit{\hat{q}}}_{C}^{2}}$, are produced as remaining unitaries. (c) Removing the $\{A,D_{1}\}$, …, $\{A,D_{m}\}$ edges by performing a series of two-mode controlled-$Z$ gates, $C_{A,D_{1}}(-1)$, …, $C_{A,D_{m}}(-1)$. (d) After performing the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state $\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{A}\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}}\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{1}}}$$\operatorname{e}^{i\operatorname{\mathit{\hat{q}}}_{B}\operatorname{\mathit{\hat{q}}}_{C}\operatorname{\mathit{\hat{q}}}_{D_{m}}}\ket{0}_{p}^{\otimes{m+3}}$ is obtained.
In Fig. 3, we couple mode $D$ with mode $A$ and use the graphical rules to show that it generates a new hyperedge of order 3 between modes $B$, $C$, and $D$. An extended example is shown in Fig. 5. We couple multiple modes $D_{1}$, …, $D_{m}$ with mode $A$ and perform the same sequence of operations which has the same effect than applying $m$ Toffoli gates on the initial state. We can see that, after performing the shearing operations on mode $A$ and a series of controlled-$Z$ gates as well as the shearing operations towards the opposite direction on modes $A$ and $D_{1}$, …, $D_{m}$ respectively, the new hypergraph state contains $m$ new hyperedges. Instead of applying a non-Gaussian Toffoli gate on a certain hypergraph state, we can apply a sequence of Gaussian operations on the same state which has the same effect.